OpenStudy (anonymous):

What is the equation of the line in slope–intercept form that passes through the point (0, 3) and is perpendicular to the line y = 2x – 6?

OpenStudy (angela210793):

perpendicular--> slope=-1/m=-1/2 y-3=-1/2(x-0) y-3=-1x/2 y=-1x/2+3